two angles are complementary. the measure of the first angle is 3x and the measure of the second angle is 7x. find the measure of each angle.
Answer by drj(1380)

Two angles are complementary. The measure of the first angle is 3x and the measure of the second angle is 7x. Find the measure of each angle.

Step 1. Complementary angles is when two angles add up to 90 degrees.

Step 2. Let 3x be the angle.

Step 3. Let 90-3x=7x be its complementary angle

Step 4. Then, 90-3x+3x=7x+3x or 90=10x. Then x=9 after dividing by 10 to both sides of the angles. Then 3x=27 degrees and 7x=63 degrees.

Step 5. ANSWER: The angles are 27 and 63 degrees.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

And good luck in your studies!

Respectfully,
Dr J
http://www.FreedomUniversity.TV